### v2

# Human-Object Interaction from Human-Level Instructions

Abstract
Intelligent agents must autonomously interact with the environments to perform daily tasks based on human-level instructions. They need a foundational understanding of the world to accurately interpret these instructions, along with precise low-level movement and interaction skills to execute the derived actions. In this work, we propose the first complete system for synthesizing physically plausible, long-horizon human-object interactions for object manipulation in contextual environments, driven by human-level instructions. We leverage large language models (LLMs) to interpret the input instructions into detailed execution plans. Unlike prior work, our system is capable of generating detailed finger-object interactions, in seamless coordination with full-body movements. We also train a policy to track generated motions in physics simulation via reinforcement learning (RL) to ensure physical plausibility of the motion. Our experiments demonstrate the effectiveness of our system in synthesizing realistic interactions with diverse objects in complex environments, highlighting its potential for real-world applications.
Figure 1: Given human-level instructions (top), we first use a diffusion-based motion generator to synthesize natural full-body human motion, finger motion, and object motion to accomplish the task (middle). Next, we employ a physics-based tracking method to imitate the generated interactions which produces physically plausible interactions (bottom). Please refer to our project page for more results.
1 Introduction
Synthesizing physically plausible, long-horizon, human-object interactions in contextual environments is crucial for applications in computer graphics, embodied AI, and robotics. When given human-level language instructions, intelligent agents should be able to comprehend these high-level commands and map them to executable action sequences. For instance, when instructed to set up a workspace, an agent must understand the abstract concept of “workspace”, identify the relevant furniture and objects in the room, and translate this concept into a series of concrete actions, such as moving and orienting a desk, placing a chair nearby, and positioning a monitor on the desk. In addition to understanding the instructions, the agent must execute actions that lead to realistic and functional human movements, including navigating cluttered spaces, manipulating various objects for the task, and appropriately releasing them after use.
Prior research has made significant progress in synthesizing human-scene interactions using large language model (LLM) planners [57]. However, most of these interactions have been limited to static objects without hand manipulation. Some studies [24] have demonstrated the manipulation of dynamic objects, but the resulting interactions still lack realistic finger movements. Recently, physics-based RL policies has demonstrated promising results in human-object interactions [31, 52]. However, generating long-horizon motions that interact with various objects in a scene-aware manner remains an ongoing challenge. Build upon these prior work, we introduce the first complete system capable of synthesizing physically accurate, long-horizon human-object interactions with synchronous full-body and intricate finger movements, from human-level language instructions.
Synthesizing coordinated full-body and finger movements from human-level instructions presents two main challenges. First, human-level instructions typically provide only a high-level task outline, requiring common sense and world knowledge to translate them into precise object arrangements and detailed execution plans. Second, existing datasets with paired full-body and finger motion data are limited to interactions with small objects and lack locomotion during object manipulation [43, 8]. Although a recent dataset [25] includes locomotion and manipulation for large objects, it lacks detailed finger movements due to the difficulty of capturing both modalities simultaneously.
To tackle the challenge of comprehending human-level instructions, we utilize the latest advancements in LLMs, which are highly effective in interpreting high-level commands[34, 28]. However, LLMs are not well-suited for directly predicting precise scene layouts, such as the exact positions and orientations of objects [1, 16]. Instead of prompting LLMs to directly specify object placements, we use object spatial relationships as intermediate representations and ask LLMs to derive these relationships. We then propose an algorithm to calculate the precise object poses based on these relationships.
For the low-level motion synthesis, we take the approach of modeling human motion from real-world human data. To address the lack of large-scale object manipulation datasets with synchronized full-body and finger motions, one could use two separate datasets to train a full-body manipulation model and another detailed finger manipulation model. However, simply fusing the two motions together will cause glaring artifacts due to the disagreement in two separately trained models. Our solution is to break this process into three steps. We first generate full-body and object motions without detailed fingers. The object motion and rough wrist poses provide guidance for generating contact-consistent finger motion in the second step. Finally, we generate full-body and object motions again conditioned on precise finger motions. This approach enables our framework to successfully produce synchronized object motion, full-body motion, and finger motions without requiring a dataset that includes all data modalities.
The generated motion, however, is not guaranteed to be physically accurate. To ensure physical plausibility, we further use reinforcement learning (RL) to track the generated motion in a physics simulator. Our method supports tracking diverse interactions with multiple distinct objects within a single sequence.
We evaluate our method through extensive quantitative and qualitative comparisons, as well as human studies, demonstrating that our approach outperforms all others. This underscores the effectiveness of our method and its potential for real-world applications.
2 Related Work
Contextual Interaction Synthesis.
Recent work on modeling human-object interactions falls into two categories: interactions with static and dynamic objects. Synthesizing human motions in static 3D scenes has been extensively studied [13, 53, 2, 21, 14, 69]. Prior research has explored regression models [48, 49, 14, 2, 32] and diffusion models [18, 21, 66, 54] to generate human motions interacting with 3D scenes, such as sitting on a chair or lying down on a sofa. Some studies have also explored reinforcement learning methods to generate physically plausible human motions [15, 71, 57]. These works focus on static objects, whereas our work focuses on dynamic interactions.
There has been increasing attention on synthesizing human-object interactions with dynamic objects [6, 35, 55, 62]. OMOMO [25] collected a dataset for dynamic human-object interactions and proposed a framework to synthesize human motions from object motions. InterDiff [61] predicts future interaction motions based on past object and human motion. CHOIS [24] synthesizes interactions in 3D scenes from text and sparse waypoints. Despite these advancements, all of these works focus on full-body human motion generation without detailed finger motions. In this work, we aim to generate synchronized object motions, human motions, and finger motions.
Hand-Object Interaction Synthesis.
Hand-object interaction synthesis has been extensively explored. In the realm of static pose generation, traditional methods often rely on physics-based control or optimization techniques to generate grasp candidates [27, 38, 41]. Some recent work leverages deep learning to directly estimate hand grasps from extensive interaction datasets [19, 3, 23, 43, 8, 5]. Beyond static grasping, another line of work also explores dynamic object manipulation [68, 65, 30].
Recent works further explore integrating whole-body motion and hand motion [45, 46, 56, 44]. IMoS [10] generates hand-object interaction using paired human motion and hand motion, while [4] leverage reinforcement learning to achieve physically plausible motion. However, these methods are constrained to manipulate small-sized objects in the GRAB dataset [43], unable to generalize to large objects. In this work, we aim to generate realistic full-body and finger motions for manipulating large objects. Since no such paired dataset exists, we leverage DexGraspNet [50] to generate realistic grasp poses, train a conditional diffusion model on GRAB [43] to generate approaching and releasing finger motions, and then integrate these with our full-body motion generation.
Physics-Based Motion Synthesis.
Physics-based methods control a simulated character to perform motion in a simulator, ensuring the physical realism of the generated motion. To make the motion more natural, many approaches leverage motion capture data to produce movements that match the style of the data [36, 37, 35, 29]. DeepMimic [36] proposes to track motion capture data using RL. Subsequent works [37, 59] leverage generative adversarial networks (GANs) to learn motion styles. Recent research also explores this approach for human-object interactions, such as playing musical instruments [60, 51] or playing soccer [58]. PhysHOI [52] learns basketball-playing skills from human demonstrations. OmniGrasp [31] enables a character to grasp and move an object along a specified object trajectory. In this work, we leverage RL to track generated kinematic motion, successfully producing long-horizon motions that interact with various objects.
LLM-based Planning.
LLMs have been widely used in reasoning, planning tasks [16, 17, 47, 67, 7] for their high-level planning capabilities. They have also been adopted in motion synthesis tasks [64, 70, 57, 62]. UniHSI [57] uses LLMs to extract action plans in a chain-of-contact format, guiding a low-level controller to synthesize corresponding motions. InterDreamer [62] employs LLMs to identify object parts that humans interact with and uses this information to retrieve initial human poses. In this work, we leverage LLMs to translate abstract human-level instructions into target scene maps and detailed execution plans, which guide the low-level interaction synthesis.
3 System Overview
Figure 2: Our system takes the scene description and human-level instruction as input and uses a high-level planner to obtain the scene map and a detailed execution plan. The low-level motion generator then generates synchronized object motion, full-body human motion, and finger motion. Finally, the physics tracker uses RL to track the generated motion, producing physically plausible motion.
Given a human-level instruction and a scene description, we develop a system that generates synchronized human motion and object motion to accomplish the task. An overview of our system is depicted in Fig. 2. The system consists of a high-level planner, a low-level motion generator, and a physics tracker. The high-level planner leverages LLMs to reason about instructions and the scene description, generating a scene map along with a detailed execution plan. The scene map specifies the desired positions and orientations of the objects, while the execution plan outlines the steps to interact with these objects. The low-level motion generator unites conditional diffusion models trained on separate datasets for full body motion and dexterous hand motion, producing synchronous object and human movements. The physics tracker then uses RL to track the generated motion in a physics simulator, ensuring physical plausibility.
4 High-Level LLM Planner
Inputs to the high-level planner consist of an initial scene description and human-level instructions. The scene description includes a list of objects in a 3D scene and their layout, while the human-level instruction describes a task that requires grounded reasoning based on world knowledge and common sense. For example, the instruction “I want to do yoga in front of the TV” implies that any obstacles in front of the TV should be moved elsewhere.
Outputs of the high-level planner comprise the scene map and the execution plan. The scene map specifies the target position $p$ and orientation $q$ for each object $o$ that requires movement, represented as $\left\{\left\{o_{1},p_{1},q_{1}\right\},\cdots,\left\{o_{n},p_{n},q_{n}\right\}\right\}$. The execution plan is a sequence of text actions associated with objects in the scene map. For example, “lift the monitor, move the monitor, put down the monitor”.
4.1 Scene Map Generation
To obtain a reliable scene map, inspired by recent works in room layout generation [1, 63], we first instruct LLMs to reason about spatial relationships of the objects as an intermediate representation from which the 3D positions and orientations of the objects are calculated subsequently. We define the following relation functions to capture spatial relationships, with flexibility to include additional relationships if needed.
  1. 1.
    on(object1, object2).
  2. 2.
    adjacent(object1, object2, direction, distance).
  3. 3.
    facing(object1, object2).
For example, if a monitor needs to be put on the table facing the chair, we can use on(monitor, table) and facing(monitor, chair) to describe its pose. If a table is positioned 1 meter to the north of the door, we can use adjacent(table, door, north, 1) to describe the relationship. We prompt LLMs to generate the spatial relationships of the objects according to the given instruction.
Given the spatial relationships, we propose an algorithm to calculate the 3D positions of each object in the scene. The orientations of objects, enforced by “facing” relations, will be addressed in the subsequent step.
The algorithm first constructs a scene graph [22] to organize the objects (nodes) and their spatial relationships (edges), which are directional, pointing from object2 to object1. The algorithm employs a method similar to topological sorting, assuming no cycles are present in the graph. The scene graph is initialized by visiting the nodes associated with static objects and setting their 3D positions as the given values. In each iteration, the algorithm looks for a new node whose 3D position has not been determined but all its predecessors have. By calculating an offset from the predecessors’ known positions, we can determine the position of the new node. For example, if object $o_{1}$ is “on” object $o_{2}$, then $o_{1}$’s height is set to match the top surface of $o_{2}$, while the horizontal position is sampled within the horizontal range of $o_{2}$. If $o_{1}$ is “adjacent” to $o_{2}$, we use the direction and distance given by the LLMs as the offset. We have included pseudo-code in the supplementary material.
To compute the orientation enforced by facing($o_{1}$, $o_{2}$), we rotate the canonical direction of $o_{1}$ in its own local frame to align with the vector from the position of $o_{1}$ to the position of $o_{2}$. The canonical direction of each object is part of the input 3D geometry information.
4.2 Execution Plan Generation
After determining the scene map, the planner must specify the interaction order with each object to ensure natural motion. For instance, if a vase is on a table, the vase should be moved before the table. LLMs are asked to generate this action sequence in a natural order and provide reasoning for it. The resulting execution plan consists of textual actions, $\left\{l_{1},l_{2},\dots,l_{T}\right\}$, where each action $l_{t}$ defines interaction involving one object. Each $l_{t}$ serves as the textual input for the low-level generator, following the format: “lift the object, move the object, put down the object”. This format is designed to closely match the textual style of the training dataset of the low-level generator.
With the scene map and execution plan in place, an A* planner generates collision-free paths to execute the plan. The path is a series of 2D waypoints in the scene, which then serve as inputs for the low-level motion generator.
5 Low-Level Motion Generator
Figure 3: The interaction module consists of four stages. Starting with the initial human and object poses, waypoints, final object pose, and a text instruction, the first stage generates coarse object and human motion. The relative wrist pose from this motion is then used to generate a valid grasp pose. Using this as a condition, the next stage re-generates the motion to align with the grasp constraint. Finally, the finger motion generation module produces finger motions during the approach and release of objects.
Given the waypoints and textual instructions generated by the high-level planner, the low-level motion generator executes an interaction module and a navigation module sequentially to generate long-horizon human-object interaction motions.
5.1 Background: Conditional Diffusion Model
The diffusion model involves a forward diffusion process and a reverse diffusion process. Starting from clean data $\bm{x}_{0}$, noise is progressively introduced until, after N steps, the data becomes approximately $\bm{x}_{N}\sim\mathcal{N}(0,\bm{I})$. The denoising network is trained to reverse this process so that sampling a noise $\bm{x}_{N}$ and applying N denoising steps yields the clean data representation $\hat{\bm{x}}_{0}$. Each denoising step is defined as
$p_{\theta}(\bm{x}_{n-1}|\bm{x}_{n},\bm{c}):=\mathcal{N}(\bm{x}_{n-1};\bm{\mu}_{\theta}(\bm{x}_{n},n,\bm{c}),\sigma_{n}^{2}\bm{I}),$ (1)
where $p_{\theta}$ represents a neural network, $\bm{c}$ represents the conditions, $\bm{\mu}_{\theta}(\bm{x}_{n},n,\bm{c})$ represents the learned mean, $\sigma_{n}^{2}$ denotes a fixed variance. The training loss is defined as $\mathcal{L}=\mathbb{E}_{\bm{x}_{0},n}||\hat{\bm{x}}_{\theta}(\bm{x}_{n},n,\bm{c})-\bm{x}_{0}||_{1}$.
5.2 Interaction Module
We introduce a multi-stage interaction module that synthesizes synchronized full-body, finger, and object motion, as shown in Fig. 3. The first stage, CoarseNet, generates initial human and object motions without detailed finger movement. Recognizing that finger poses typically remain static during interaction, the second stage generates a grasp pose based on the initial results. Then RefineNet re-generates the human and object motions to seamlessly align with the optimized grasp pose. Finally, FingerNet generates smooth finger motions for approaching and releasing the object, completing the interaction sequence.
5.2.1 CoarseNet
Given the initial human and object poses, waypoints, and final object pose, along with a text instruction, CoarseNet generates synchronized human motion and object motion. This is achieved using a pre-trained diffusion model [24].
Data Representation.
We denote human motion as $\bm{H}\in\mathbb{R}^{T\times D}$, where $T$ represents the sequence length and $D$ represents the dimension of the human pose. Each $\bm{H}_{t}$ includes global joint positions and 6D rotations [72], excluding finger joints. Object motion $\bm{O}\in\mathbb{R}^{T\times 12}$ includes the object’s global position and rotation matrix. CoarseNet predicts human motion, object motion, and contact labels $\bm{L}\in\mathbb{R}^{2}$ for both hands, resulting in $\bm{x}=\left\{\bm{H},\bm{O},\bm{L}\right\}$.
The input condition is represented as $\bm{c}=\left\{\bm{S},\bm{G},\bm{T}\right\}$. $\bm{S}\in\mathbb{R}^{T\times(D+12)}$ is a masked motion data representation. It includes the human and object pose in the first frame, the object pose in the last frame, and 2D waypoints sampled every 30 frames. Any remaining entries in $\bm{S}$ are padded with zeros. The object geometry $\bm{G}$ represents the BPS representation [39] to encode object geometry, and $\bm{T}$ refers to the text embeddings extracted using CLIP [40].
Segmenting Contact Phases.
We segment the generated interaction sequence into pre-contact, contact, and post-contact phases based on the predicted contact labels $\bm{L}$. The contact phase is determined separately for each hand, dividing the model output $\bm{x}_{\text{coarse}}$ into $\bm{x}_{\text{pre-contact}}\oplus\bm{x}_{\text{contact}}\oplus\bm{x}_{\text{post-contact}}$. We then compute the average wrist pose $\bm{w}$ from $\bm{x}_{\text{contact}}$, serving as input for the next stage.
5.2.2 Grasp Pose Generation
The second stage generates the finger pose for the contact phase using a state-of-the-art grasp pose generation method [50]. Given an object, wrist pose $\bm{w}$ and rest finger pose, an optimization is conducted to generate a physically plausible grasp pose. We modified [50] to fit our needs by: (1) increasing the object surface sample points from 2,000 to 20,000 to maintain sample density for larger objects and (2) omitting the force closure term for tasks involving both hands, such as carrying a box, where stability doesn’t solely rely on a single hand.
The optimized grasp pose is denoted by $\hat{\bm{g}}=(\hat{\bm{w}},{\hat{\bm{\theta}}})$, where $\hat{\bm{w}}$ is the optimized wrist pose, and $\hat{\bm{\theta}}$ is the finger pose. This grasp pose $\hat{\bm{g}}$ is maintained throughout the contact phase. However, as CoarseNet’s generated motion may not fully align with $\hat{\bm{w}}$, we introduce RefineNet to address this misalignment.
5.2.3 RefineNet
The goal of RefineNet is to re-generate human-object interaction motions that align with the optimized grasp pose $\hat{\bm{g}}$. We adopt a conditional diffusion model based on CoarseNet, adding two conditions: wrist-object relative pose to align the wrist with $\hat{\bm{w}}$ and object static to keep the object stationary during non-contact phases. RefineNet’s input condition, $\bm{c}_{r}=\left\{\bm{W},\bm{S}_{r},\bm{G},\bm{T}\right\}$, includes $\bm{G}$ and $\bm{T}$ from CoarseNet, with $\bm{W}$ and $\bm{S}_{r}$ detailed as follows.
Wrist-Object Relative Pose Condition.
The condition $\bm{W}\in\mathbb{R}^{T\times 18}$ represents the position and 6D rotation of both wrists in the object’s frame, including only contact-phase wrist poses, with other entries set to zero.
Object Static Condition.
In CoarseNet, the masked motion condition $\bm{S}$ was zero-padded except at the start, end, and waypoints. RefineNet uses contact phase information from CoarseNet to adjust $\bm{S}$ to $\bm{S_{r}}$, assigning a static object pose in pre-contact and post-contact frames.
Wrist-Object Relative Pose Loss.
We incorporate an additional loss to enforce the wrist-object relative pose condition. We sample 100 points $\bm{K}_{\text{rest}}\in\mathbb{R}^{100\times 3}$ uniformly on the rest object’s surface and pre-calculate their positions in both wrists’ coordinate frames, denoted as $\bm{K}_{\text{w}}\in\mathbb{R}^{2\times T\times 100\times 3}$. At each time step, we compute the points’ positions in the wrist frame, $\hat{\bm{K}}_{\text{w}}$, using the predicted orientations and positions of the object and wrist, $\bm{R}_{\text{o}},\bm{T}_{\text{o}},\bm{R}_{\text{w}},\bm{T}_{\text{w}}$:
$\bm{K}_{\text{global}}$ $=\bm{R}_{\text{o}}\bm{K}_{\text{rest}}+\bm{T}_{\text{o}},$ (2)
$\hat{\bm{K}}_{\text{w}}$ $=\bm{R}_{\text{w}}^{-1}\left(\bm{K}_{\text{global}}-\bm{T}_{\text{w}}\right).$ (3)
The loss is computed as:
$\mathcal{L}_{\text{relative }}=\sum_{t=1}^{T}\bm{L}_{t}\left\|\hat{\bm{K}}_{\text{w},t}-\bm{K}_{\text{w},t}\right\|_{1},$ (4)
where $\bm{L}_{t}$ represents the contact labels, masking out the loss when the hand and object are not in contact.
Post-Processing.
RefineNet generates human-object motions that align closely with input conditions, though minor adjustments may be required. To refine alignment, we apply post-processing: replacing object poses in pre- and post-contact phases with static poses and recalculating wrist trajectories during the contact phase based on object motion and the optimized wrist pose $\hat{\bm{w}}$. Additional implementation details are in the supplemental materials.
5.2.4 FingerNet
In this stage, we generate finger motions for the pre-contact and post-contact phases to create natural transitions between the rest and grasp poses, forming a complete interaction sequence along with predictions in prior stages. Given the start and end finger poses and the wrist trajectory in between, we employ a conditional diffusion model to predict the finger motions denoted as $\bm{F}\in\mathbb{R}^{T\times D^{\prime}}$, representing local 6D rotations [72].
The model is conditioned on $\bm{c}_{f}=\left\{\bm{P},\bm{F}_{s},\bm{F}_{e}\right\}$, where $\bm{P}$ captures spatial hand-object relationships and $\bm{F}_{s},\bm{F}_{e}$ are the start and end finger poses. Following prior work [68, 45], $\bm{P}\in\mathbb{R}^{T\times 100}$ is computed by sampling 100 palm-side mesh vertices and measuring their closest distances to the object, with the finger pose set to mean pose when calculating $\bm{P}$.
5.3 Navigation Module
To generate long sequences of interactions with multiple objects, the human needs to navigate through the scene based on waypoints. We adopt a conditional diffusion model to generate human navigation motion $\bm{H}\in\mathbb{R}^{T\times D}$. The input conditions consist of the initial human pose, 2D waypoints, and waypoint orientations (expressed as normalized direction vectors) to maintain consistent facing directions. Smooth transitions between the interaction and navigation modules are achieved by using the final pose of one module as the starting pose for the next.
6 Physics Tracker
After obtaining a sequence of kinematic motions, we train a tracking policy using reinforcement learning to control a physically simulated character that imitates these motions. This approach helps eliminate artifacts (e.g., hand-object penetration and foot sliding) in the kinematic motions and ensures the physical plausibility of the generated results.
We use PPO [42] as the backbone reinforcement learning algorithm, and adopt an importance sampling strategy to facilitate the tracking of long motion sequences involving interactions with multiple objects. Policies trained by our approach can accurately track the kinematics motions, enabling the character to grasp and move distinct objects along the generated trajectories. Implementation and training details for the physics tracker are provided in the supplementary materials.
7 Experiment
We evaluate each system module by comparing our proposed approach with baseline methods and conducting ablation studies. We highly recommend that readers refer to the supplementary video for additional qualitative results.
7.1 Evaluation of High-Level Planner
The core component of the high-level planner is the scene map calculation. To assess its effectiveness, we compare our method to a baseline that uses LLMs to directly predict object positions and orientations. We design 15 human-level instructions covering common everyday tasks for assessment, using OpenAI’s GPT-4o [33] for both methods.
Figure 4: Comparison between (a) the baseline and (b) our LLM planner. The baseline generates an incorrect monitor position and orientation, and places the chair in a position that intersects with the table. The human perception study results (bottom) show that the majority of participants preferred our results.
[…]
Table 1: Interation synthesis on the FullBodyManipulation dataset [25]. Our full model achieves superior interaction quality compared to all others.
Figure 5: Qualitative comparisons of each method. All baselines and ablations cannot produce precise hand and finger motions, while our method generates natural hand-object interactions.
Figure 6: Comparison of our full model and baseline models through human perception studies.
We conduct a human perception study to evaluate the methods, with results presented in Fig. 4. Each pair was reviewed by 20 Amazon Mechanical Turk (AMT) workers. As shown in Fig. 4, the baseline method fails to correctly position the objects and accurately determine their facing directions.
7.2 Evaluation of Low-Level Motion Generator
Datasets.
CoarseNet and RefineNet are trained on the FullBodyManipulation dataset [25], which includes 10 hours of human-object interaction with 15 objects but lacks finger motion. FingerNet uses the GRAB dataset [43], containing full-body and finger motion for 10 subjects with 51 objects. The navigation module is trained on HumanML3D [12], featuring 28 hours of diverse motions with language descriptions. For all datasets, we follow standard data partitioning from prior work [25, 43, 12].
Evaluation Metrics.
Following prior work [24], we evaluate our approach from various aspects. For condition matching, we report the waypoint following errors ($T_{xy}$) in centimeters. For human motion quality, we report the foot sliding score (FS) and foot height ($H_{\text{feet}}$), both measured in centimeters. For interaction quality, we use contact precision (Cprec), recall (Crec), and F1 score (CF1). To better assess finger-object interaction, we also report contact coverage (CC) – the percentage of hand points within -2mm to +2mm of the object surface – and intersection volume (IV), the overlap volume of hand and object meshes in $\text{cm}^{3}$, following [11]. For ground truth difference, we report the mean per-joint position error (MPJPE), object position error ($T_{\text{obj}}$), both in centimeters (cm), along with the object orientation error ($O_{\text{obj}}$). We also conduct a human perception study to measure the motion quality.
Baselines.
For the task of generating full-body human motion, finger motion, and object motion from text, prior work [9, 26] trained models on GRAB [43] to predict motions that manipulate given objects. GRAB [43] contains small object manipulation data without locomotion. Models trained on GRAB cannot generalize to scenarios involving large-sized objects that require full-body coordination and locomotion. The approach from [9, 26] cannot be directly applied to train on the FullBodyManipulation dataset [25] since this dataset does not contain finger motions, making it infeasible to learn a model via direct supervision. Thus, we establish our baseline by combining two complementary methods: CHOIS [24], which generates object motion and full-body motion from text, and GRIP [45], which refines arm movements and generates finger motions based on CHOIS’s output.
In addition, to evaluate the impact of each component, we compare the full system to two ablations: (1) CoarseNet alone (CNet) and (2) CoarseNet + RefineNet (C+RNet). This evaluation focuses on the motion synthesis quality before physics tracking. We present the evaluation for the physics tracker in the next subsection.
Results.
The qualitative results are shown in Fig. 5. Both ablations exhibit a lack of finger details, and GRIP [45] fails to generate accurate hand motion, leading to significant artifacts. In contrast, our method generates precise finger-object interactions. The quantitative results in Tab. 1 demonstrate our method’s superior performance in terms of interaction quality. We also observe the impact of RefineNet: compared to CNet, C+RNet achieves better interaction quality. As shown in Fig. 5, C+RNet maintains reasonable relative poses between the hand and object throughout the interaction, while CNet exhibits more artifacts.
We further conduct a human perception study to complement our evaluation. We generate 15 sequences using each method and compose 60 pairs, with each pair consisting of our results and the results of a baseline. Each pair was reviewed by 20 Amazon Mechanical Turk (AMT) workers. As shown in Fig. 6, our method is consistently preferred over the baselines by a large margin. Notably, our method achieves even higher preference rates when compared to the real-world data from the FullBodyManipulation [25] dataset, as we generate detailed finger motions that are absent in the original dataset.
7.3 Evaluation of Physics Tracker
The physics tracker can accurately track kinematic motion with an average tracking error of approximately 2 cm per frame per joint. Tracking error is defined as the position error between the reference motion and the tracked motion. Through physics simulation, artifacts such as foot floating and human-object penetration are eliminated, and foot sliding is also significantly reduced, with an FS metric of 0.002. Qualitative results are shown in Fig. 7.
We also compare our approach with PhysHOI [52]. While [52] demonstrates impressive results in basketball playing motion, it does not perform well in our task. The policy fails to grasp objects, as shown in Fig. 8.
Figure 7: Qualitative comparison of kinematic motion and physics-based motion. Our physics tracker corrects artifacts such as foot floating and penetration.
Figure 8: PhysHOI [52] fails to track kinematic motion, while our policy successfully drives the character to interact with diverse objects.
8 Conclusions
In this work, we introduced a system that simulates realistic human-object manipulation in a contextual environment based on human-level language instructions. The framework consists of a high-level LLM planner that infers a target scene layout and execution plan from the given instructions, a low-level motion generator that generates synchronized, full-body, fingers, and object motion, and a physics tracker that imitate the generated kinematic motions and produces physically plausible interaction motions. Our system effectively handles a variety of tasks, showcasing its potential for various real-world applications.
Limitations.
Our system currently uses text or image inputs to represent the contextual environment, and could potentially be enhanced with more structured representations like voxel grids as used in [20]. While our approach successfully leverages LLMs to interpret human-level instructions, there is room to improve plan generation for more complex environments. For grasp pose generation, we build upon [50], though alternative approaches could be explored to further improve grasp success rates.
Future Work.
A promising direction is creating an agent with egocentric vision perception as input to LLM/VLMs for high-level motion planning. This could enable the agent to exhibit more natural behaviors, such as purposeful head orientation for observation and better body positioning to avoid collisions.
Acknowledgement.
We thank Wenyang Zhou for data processing. This work is in part supported by the Wu Tsai Human Performance Alliance at Stanford University and the Stanford Institute for Human-Centered AI (HAI).